OpenStudy (anonymous):

after arnold pumps gas into the gas tank of his truck, the gas tank foes from 1/8 full to 7/8 full. if the gas costs $4.75 per gallon and Arnold is charged 57.00 for gas, what is the capacity, in gallons, of the gas tank?

OpenStudy (anonymous):

16 gallons is the capacity

OpenStudy (anonymous):

work please?

OpenStudy (anonymous):

57.00 / 4.75 = 12

OpenStudy (anonymous):

(3/4)x = 12 (4/3)(3/4)x = (4/3)12 1x = 48/3 x = 16

OpenStudy (anonymous):

7/8 - 1/8 is 6/8 which is 3/4